OpenStudy (anonymous):

Trapezoid with a short base of 2 meters, a long base of 10 meters and a height of 3 meters whats the area

OpenStudy (waheguru):

http://math.about.com/od/formulas/ss/areaperimeter_2.htm the formula

OpenStudy (angela210793):

A=(B+b)*h/2=(2+10)*3/2=18m^2

OpenStudy (waheguru):

thx now i ama life saver